OpenStudy (anonymous):

Given the arithmetic sequence an = 2 − 5(n + 1), what is the domain for n?

OpenStudy (anonymous):

All real numbers All integers where n ≥ 1 All integers where n > 1 All integers where n ≥ 2

OpenStudy (anonymous):

n sum k k=1

OpenStudy (anonymous):

= (n2 + n) / 2 = (1/2)n2 + (1/2)n

OpenStudy (anonymous):

so all real numbers?

OpenStudy (anonymous):

yes all r real number.